Having a baby in a U.S. hospital costs $2528, on average. A study of 200 AML (Active Management of Labor) deliveries had a mean cost of $2480 with a standard deviation of $766. At the 5% significance level, do the data provide sufficient evidence to conclude, that on average, AML reduces the cost of having a baby in a U.S. hospital? Perform the hypothesis test and provide the results.

A. Do not reject Ho

B. Do not reject Ha

C. No results

D. Reject Ho
Answer #1

Solution:

This is a left (One) tailed test,

The null and alternative hypothesis is,  

Ho: 2528

Ha: 2528

The test statistics,

t = ( - )/ (s/)

= ( 2480 - 2528 ) / ( 766 / 200 )

= -0.886

df = n - 1 = 199

P-value = 0.1883

Since, P-value > level of significance, Do not reject H0.

Do not reject Ho

Option A is correct.